OpenStudy (anonymous):

Find the products. 2 1/4 x 11 5/6 = 3 2/3 x 10 1/2 = 2 1/2 x 1 7/10 = 6 x 3 3/8 = Please explain.

OpenStudy (anonymous):

change them into improper fraction first. use the first as example: 2 1/4 x 11 5/6 =9/4 x 71/6 then, denominator times denominator, numerator times numerator. =(9x 71)/(4x 6) =639/24 then, you simplify it. In this case, both denominator and numerator can be divided by 3 =213/8 you can try the other question and post your answer to see if your answer is correct or not.

OpenStudy (blues):

You can also try posting in the correct group ~ math, not biology.

OpenStudy (anonymous):

Sorry!